OpenStudy (anonymous):

Evaluate the funtion fo the given value of x f(x)=4x+1 for x=3

OpenStudy (anonymous):

Evaluate the funtion fo the given value of f(x)=2x-11 for x=5

OpenStudy (anonymous):

Evaluate the funtion fo the given value of \[f(x)=\frac{ 5 }{ 2 }x+8 for x=4 \]

OpenStudy (anonymous):

f(3)=4(3)+1=12+1=13 this means for the function f(x) sub x=3 so replace al x's with 3 and solve

OpenStudy (anonymous):

1) f(3) = 4(3) + 1 3f = 12 + 1 3f = 13 f = 13/3

OpenStudy (anonymous):

2) f(5) = 2(5) - 11 5f = 10 - 11 5f = -1 f = -1/5

OpenStudy (anonymous):

do the same for the other questions and tell me what you get

OpenStudy (anonymous):

PINAY17 thats not correct

OpenStudy (anonymous):

3) f(4) = 5/2 (4) +8 4f = 10 + 8 4f = 18 f = 9/2

OpenStudy (anonymous):

PINAY17 those are all wrong, sorry